A 400 g mass A trolley moving with velocity of 2.0ms-¹ collides with a 500 g mass B trolley that moves in opposite direction with a velocity of 2.5ms-¹. If the cart B moves with a velocity of 1.5ms-¹ in the same direction after the collison, what is the velocity of the trolley A after the collison ?


1
Expert's answer
2021-04-13T06:36:52-0400

We can find the velocity of the trolley A after the collision from the law of conservation of momentum:


"m_Av_{Ai}-m_Bv_{Bi}=m_Av_{Af}-m_Bv_{Bf},""v_{Af}=\dfrac{m_Av_{Ai}-m_Bv_{Bi}+m_Bv_{Bf}}{m_A},""v_{Af}=\dfrac{0.4\ kg\cdot2\ \dfrac{m}{s}-0.5\ kg\cdot2.5\ \dfrac{m}{s}+0.5\ kg\cdot1.5\ \dfrac{m}{s}}{0.4\ kg},""v_{Af}=0.3\ \dfrac{m}{s}."